OpenStudy (anonymous):

Find the dimensions of the rectangle with area 180 square inches and smallest possible perimeter.

OpenStudy (anonymous):

the two variables are x and z. how do I set this up in my graphing calculator to get the best answer?

OpenStudy (anonymous):

A=x*y =180 ===> P =2(x+y) ===> P=2(x+180/x) ===> P' =2-360/x^2 ===> P' = 0 ===> x=6sqrt(5) ==> y = 180/6sqrt(5)

OpenStudy (anonymous):



OpenStudy (anonymous):

hence. from the graph at x= 6sqrt(5) is the smallest possible perimeter.

OpenStudy (anonymous):

Make factor pairs of 1801*180 1*180 2*90 3*60 4*45 5*36 6*30 9*20 10*18 12*15

OpenStudy (anonymous):

Smallest perimeter means a square, so the side is sqrt of 180. \[\sqrt{180} = \sqrt{9*4*5} = 6\sqrt{5}\]

OpenStudy (anonymous):

Try to find perimeter with each of these factor pair 1*180 ----- p=362 2*90 -------- p=184 3*60 126 4*45 98 5*36 82 6*30 72 9*20 58 10*18 56 12*15 54 As you can see, smallest possible perimeter is possible with dimensions as 12 and 15

OpenStudy (anonymous):

okay, I see that. and if I wanted to put this as y= in my graphing calculator to sketch the graph how would I do that?